# Finitary conditions in special semidirect products

### v1

@@ 4.3. The finitely equated conditions
Suppose that $M$ is finitely right equated. Given a finite generating set $X$ for $\mathbf{r}(a)$, setting
$Y:=\{((\emptyset,1),(a^{-1}A,1)),((a^{-1}A,1),(\emptyset,1)),((a^{-1}A,p),(a^{-1}A,q)):(p,q)\in X\},$
is sufficient to generate those elements $((U,u),(V,v))$ of $\mathbf{r}((A,a))$ satisfying $U,V\subseteq a^{-1}A$. Indeed, if $u=p_{1}t_{1},q_{1}t_{1}=p_{2}t_{2},\cdots q_{n}t_{n}=v$ is an $X$-sequence for $(u,v)$, then a $Y$-sequence for $((U,u),(V,v))$ is:
$(U,u)=(\emptyset,1)(U,u),\,\,(a^{-1}A,1)(U,u)$ $=$ $(a^{-1}A,p_{1})(\emptyset,t_{1}),$
$(a^{-1}A,q_{1})(\emptyset,t_{1})$ $=$ $(a^{-1}A,p_{2})(\emptyset,t_{2})$
$\vdots$
$(a^{-1}A,q_{n})(\emptyset,t_{n})$ $=$ $(a^{-1}A,1)(V,v),\,\,(\emptyset,1)(V,v)=(V,v).$
In particular, this demonstrates that each right annihilator of the form $\mathbf{r}((M,a))$ is finitely generated. For an arbitrary subset $A\subseteq M$, the challenge then is finite generation of pairs $((U,u),(V,v))$ where $aU\setminus A=aV\setminus A\neq\emptyset$. We show that this is not possible in complete generality, and give a sufficient condition under which the result holds.
@@
The following lemma will be used show that there exist finitely right equated monoids $M$ such that $\mathcal{S}(M)$ is not finitely right equated.
@@
It can be deduced from Proposition 2.11 and Proposition 3.5 that if $M$ is right abundant, then $\mathcal{S}(M)$ is finitely right equated. This also follows from a short direct argument:
Lemma 4.18.
Let $a,e\in M$ with $a\,\mathcal{L}^{*}\,e=e^{2}$. Then for any subset $A\subseteq M$, the right annihilator congruence $\mathbf{r}((A,a))$ is finitely generated.
@@
Note that, since $ae=a$ we have $(A,a)(\emptyset,1)=(A,a)=(A,a)(a^{-1}A,e)$. Let $\rho$ denote the right congruence generated by the pair $((\emptyset,1),(a^{-1}A,e))$; we shall demonstrate that $\rho$ is equal to $\mathbf{r}((A,a))$. If $((U,u),(V,v))\in\mathbf{r}((A,a))$, then we have $A\cup aU=A\cup aV$ and $au=av$. Moreover, if $x\in eU$ and $x\not\in a^{-1}A$, then $ax\in aU\setminus A=aV\setminus A\subseteq aV=aeV$, but then $ax=aev^{\prime}$ for some $v^{\prime}\in V$ and since $a\,\mathcal{L}^{*}\,e$ we obtain $ex=ev^{\prime}$. Since $x\in eU$ note that $ex=x$ and so $eU\subseteq eV$, and by a dual argument we deduce that $eU=eV$. It follows that
$(U,u)=(\emptyset,1)(U,u)\,\rho\,(a^{-1}A,e)(U,u)$ $=$ $(a^{-1}A\cup eU,eu)$
$=$ $(a^{-1}A\cup eV,ev)=(a^{-1}A,e)(V,v)\,\rho\,(\emptyset,1)(V,v)=(V,v).$
Corollary 4.19.
Let $M$ be a right abundant monoid. Then $\mathcal{S}(M)$ is weakly right coherent if and only if $M$ is strongly right ideal Howson.
On the other hand, as we have alluded to earlier, there exist left abundant monoids $M$ such that $\mathcal{S}(M)$ is not left abundant. Analogously to Lemma 4.16 we now give a criterion for a monoid $M$ that prevents $\mathcal{S}(M)$ from being finitely left equated. We then exhibit an example of a semilattice (which is certainly left abundant) satisfying this condition (so that $\mathcal{S}(M)$ is not finitely left equated and so certainly not left abundant).
In what follows where we say that $b$ is indecomposable we mean that the only factors of $b$ are $1$ and $b$.
Lemma 4.20.
Let $M$ be a monoid and let $a\in M$ be such that there are infinitely many distinct pairs of elements $(b,c)\in\mathbf{l}(a)$ where $b$ is indecomposable and $c\not\leq_{\mathcal{R}}b$. Then $\mathbf{l}((\{1\},a))$ is not finitely generated.
Proof.
We begin by observing that for $((P,p),(Q,q))\in X$ we have that $(P,p)(\{1\},a)=(Q,q)(\{1\},a)$ and so $P\cup\{p\}=Q\cup\{q\}$ and $pa=qa$. Suppose for contradiction that $\mathbf{l}((\{1\},a))$ is generated by a finite set $X$, where we may assume that $X$ is symmetric. Choose $(b,c)\in\mathbf{l}(a)$ such that $b$ is indecomposable, $c\not\leq_{\mathcal{R}}b$ and for all $((P,p),(Q,q))\in X$ we have $P\not\subseteq\{1,b,c\}$ and $p\neq b$.
Let $U\subseteq\{1,b,c\}$ and suppose that $(U,b)=(T,t)(P,p)$ where $(T,t)\in\mathcal{S}(M)$ and there is a pair $((P,p),(Q,q))\in X$. We claim that $(T,t)(Q,q)=(V,b)$ where $V\subseteq\{1,b,c\}$. From $(U,b)=(T,t)(P,p)$ we have that $U=T\cup tP$ and $b=tp$. Since $b$ is indecomposable and $p\neq b$ we deduce that $p=1$ and $b=t$. Further, $T\cup bP\subseteq\{1,b,c\}$. Let $r\in P$. Then $br=1,b$ or $c$. The only possibility (since $c\not\leq_{\mathcal{R}}b$) is for $br=b$ so that $r\in\{1,b\}$. We now split the argument into two cases.
If $b$ is not idempotent, then $r=1$ and so $P\subseteq\{1\}$. Since $P\cup\{p\}=Q\cup\{q\}$ we have that $\{1\}=Q\cup\{q\}$ so that $Q\subseteq\{1\}$ and $q=1$. Then $(T,t)(Q,q)=(T,b)(Q,1)=(V,b)$ where $V\subseteq\{1,b,c\}$ as required.
If $b$ is idempotent, then $r\in\{1,b\}$ and so $P\subseteq\{1,b\}$. Since $P\cup\{p\}=Q\cup\{q\}$ we have that $Q\subseteq\{1,b\}$ and $q\in\{1,b\}$. Then, using the fact that $b$ is idempotent, $(T,t)(Q,q)=(T,b)(Q,q)=(V,b)$, where $V\subseteq\{1,b,c\}$ as required. This verifies our claim.
Notice now that $((\{b,c\},b),(\{b,c\},c)\in\mathbf{l}((\{1\},a))$ and as $b\neq c$ there must be an $X$-sequence of positive length connecting $(\{b,c\},b)$ to $(\{b,c\},c)$. But this is now shown to be impossible. We deduce that $\mathbf{l}((\{1\},a))$ is not finitely generated. ∎
Example 4.21.
Let $M$ be the monoid with presentation as a monoid semilattice given by
$M:=\langle a,b_{i},c_{i},i\in\mathbb{N}\mid b_{i}a=c_{i}a\rangle.$
Since $M$ is a semilattice, it is certainly abundant and so finitely left equated. By construction, we have infinitely many distinct pairs $(b_{i},c_{i})\in\mathbf{l}(a)$. It is easily seen that $b_{i}$ (indeed also $c_{i}$ and $a$) is irreducible for all $i\in\mathbb{N}$. In particular, $c_{i}$ is not $\mathcal{R}$-below $b_{i}$. Thus by Lemma 4.20 we see that $M$ is not finitely left equated.
We can build on this example to construct further examples of finitely left equated monoids satisfying the conditions of Lemma 4.20. Take $M$ as above and any cancellative monoid $T$ with trivial group of units (for example, a free monoid). By [4, Corollary 4.26] we have that $M\times T$ is finitely left equated. It is then easy to see that $(a,1_{T})$ satisfies the conditions of Lemma 4.20 so that $\mathcal{S}(M\times T)$ is not finitely left equated.
Whilst left abundance of $M$ is not sufficient to guarantee that $\mathcal{S}(M)$ is finitely left equated, we do have the following sufficient condition.
Proposition 4.22.
Let $M$ be a right cancellative monoid. Then $\mathcal{S}(M)$ is finitely left equated. Consequently, $\mathcal{S}(M)$ is weakly left coherent if and only if $M$ is left ideal Howson.
Proof.
Suppose $M$ is right cancellative. For each $\alpha=(A,a)\in\mathcal{S}(M)$, we show that the left congruence $\mathbf{l}(\alpha)=\lambda_{W}$ where $W=\{(A,1),(\emptyset,1))\}$. Certainly $W\subseteq\mathbf{l}(\alpha)$, since:
$(A,1)(A,a)=(A\cup A,a)=(A,a)=(\emptyset\cup A,a)=(\emptyset,1)(A,a).$
Notice also that for any subsets $R,T$ of $A$ we have $(T,1)\,\lambda_{W}\,(R,1)$ since:
$(T,1)=(T,1)(\emptyset,1)\,\lambda_{W}\,(T,1)(A,1)=(A,1)=(R,1)(A,1)\,\lambda_{W}\,(R,1)(\emptyset,1)=(R,1).$
Now for $((U,u),(V,v))\in\mathbf{l}(\alpha)$, by definition, we have
$(U\cup uA,ua)=(U,u)(A,a)=(V,v)(A,a)=(V\cup vA,va),$
and by right cancellativity of $M$ it follows that $u=v$ and $U\cup uA=V\cup uA$. Let $U^{\prime}=U\setminus(U\cap uA)$ and $V^{\prime}=V\setminus(V\cap uA)$ so that $U^{\prime}=V^{\prime}$ and let $R,T\subseteq A$ with $uR=U\cap uA$ and $uT=V\cap uA$. Then
$(U,u)=(U^{\prime},u)(R,1)\,\lambda_{W}\,(U^{\prime},u)(T,1)=(V^{\prime},v)(T,1)=(V,v).$
Thus for $M$ right cancellative we have that $\mathcal{S}(M)$ is weakly left coherent if and only if it is left ideal Howson, which by Proposition 4.12 is equivalent to $M$ being left co-ordinated. Since $M$ is right cancellative Proposition 4.9 gives that this is equivalent to the condition that $M$ is left ideal Howson. ∎
Corollary 4.23.
If $M$ is left (right) LCM, then $\mathcal{S}(M)$ is weakly left (right) coherent.
Proof.
The result follows directly from Proposition 4.22 and Corollary 4.19. ∎
@@
For a right congruence $\rho_{W}$ (respectively, a left congruence $\lambda_{W}$) on a monoid $M$, we may call the sequence $c_{1},d_{1},\cdots,c_{n}d_{n}$ appearing in the (left-right dual of the) $W$-sequence in Proposition 2.1 the skeleton of the $W$-sequence. If $M$ is right abundant, then we have shown in Lemma 4.18 that $\mathcal{S}(M)$ is finitely right equated and, moreover, for any fixed $(A,a)\in\mathcal{S}(M)$ we may choose a finite (indeed, singleton) generating set $W$ for $\mathbf{r}((A,a))$ such that any two elements of $\mathbf{r}((A,a))$ are connected by a $W$-sequence of length 2 and with the same skeleton. Likewise, if $M$ is right cancellative, then Proposition 4.22 yields that $\mathcal{S}(M)$ is finitely left equated and, moreover, for any fixed $(A,a)\in\mathcal{S}(M)$ we may choose a finite (indeed, singleton) generating set $W$ for $\mathbf{l}((A,a))$ such that any two elements of $\mathbf{l}((A,a))$ are connected by a $W$-sequence of length 2 and with the same skeleton. On the other hand, if condition (L) holds for $\mathcal{S}(M)$ so that in particular $\mathbf{l}((A,a))$ is finitely generated as a left $\mathcal{S}(M)$-act by a subset $W$, then any two elements of $\mathbf{l}((A,a))$ are connected by a $W$-sequence of length 1. This leads us to the following questions.
Question 4.25.
Let $M$ be a monoid. What conditions must $M$ satisfy such that for any $(A,a)\in\mathcal{S}(M)$ we have that $\mathbf{l}((A,a))$ is finitely generated by a set $W$ such that (a) only one skeleton is required (b) there is a bound on the length of $W$-sequences?

### v2

@@ 4.3. The finitely equated conditions
Conversely, suppose that $M$ is strongly finitely right equated. Fix $a\in M$ and suppose that $X$ is a finite generating set for $\mathbf{r}(a)$ with fixed skeleton $p_{1},q_{1},\ldots,p_{n},q_{n}$. For an arbitrary $A\subseteq M$ we aim to show that $\mathbf{r}((A,a))$ has the following fixed skeleton:
$(\emptyset,1),(a^{-1}A,1),(a^{-1}A,1),(\emptyset,1),(\emptyset,p_{1}),(\emptyset,q_{1}),\ldots,(\emptyset,p_{n}),(\emptyset,q_{n}),(\emptyset,1),(a^{-1}A,1),(a^{-1}A,1),(\emptyset,1).$
As a first step, observe that
$(A,a)(\emptyset,1)=(A,a)=(A,a)(a^{-1}A,1)$
$(A,a)(\emptyset,p_{i})=(A,ap_{i})=(A,aq_{i})=(A,a)(\emptyset,q_{i}),$
so that the proposed sequence indeed splits into pairs of consecutive elements belonging to $\mathbf{r}((A,a))$.
Let $((U,u),(V,v))\in\mathbf{r}((A,a))$. Then $A\cup aU=A\cup aV$ and $au=av$. Write $U=U_{1}\cup U_{2}$ and $V=V_{1}\cup V_{2}$ where $U_{1}=U\setminus a^{-1}A$, $U_{2}=U\cap a^{-1}A$, $V_{1}=V\setminus a^{-1}A$ and $V_{2}=V\cap a^{-1}A$. Then the condition $A\cup aU=A\cup aV$ becomes $A\cup aU_{1}=A\cup aV_{1}$, and since the union is disjoint we deduce that $aU_{1}=aV_{1}$, and so in particular $U_{1}=\emptyset$ if and only if $V_{1}=\emptyset$. Also note that using the first two pairs in our proposed skeleton we obtain:
$(U,u)=(\emptyset,1)(U,u),\qquad(a^{-1}A,1)(U,u)=(a^{-1}A,1)(U_{1},u),\qquad(\emptyset,1)(U_{1},u)=(U_{1},u)$
whilst using the last two pairs of the proposed skeleton we obtain:
$(V_{1},v)=(\emptyset,1)(V_{1},v),\qquad(a^{-1}A,1)(V_{1},v)=(a^{-1}A,1)(V,v),\qquad(\emptyset,1)(V,v)=(V,v).$
Thus it suffices to show that there exist elements $(T_{i},t_{i})\in\mathcal{S}(M)$ such that
$(U_{1},u)=(\emptyset,p_{1})(T_{1},t_{1}),(\emptyset,q_{1})(T_{1},t_{1})=(\emptyset,p_{2})(T_{2},t_{2}),\ldots,(\emptyset,q_{n})(T_{n},t_{n})=(V_{1},v).$
Equivalently, we need subsets $T_{i}\subseteq M$ with
$U_{1}=p_{1}T_{1},q_{1}T_{1}=p_{2}T_{2},\ldots,q_{n}T_{n}=V_{1}$
and elements $t_{i}\in M$ such that $u=p_{1}t_{1},q_{1}t_{1}=p_{2}t_{2},\ldots,q_{n}t_{n}$. The latter is secured since $M$ is strongly finitely right equated with the given skeleton. We focus then on constructing the subsets $T_{i}$.
Consider
$B:=\{(h,k):h\in U_{1},k\in V_{1},ah=ak\}\subseteq\mathbf{r}(a).$
For all $(h,k)\in B$ there exist $t_{1}(h,k),\ldots,t_{n}(h,k)\in M$ such that
$h=p_{1}t_{1}(h,k),\,q_{1}t_{1}(h,k)=p_{2}t_{2}(h,k),\,\ldots,\,q_{n}t_{n}(h,k)=k.$
Setting $T_{i}=\{t_{i}(h,k):(h,k)\in B\}$ for $i=1,\ldots,n$ we claim that $U_{1}=p_{1}T_{1},q_{1}T_{1}=p_{2}T_{2},\ldots,q_{n}T_{n}=V_{1}$, as required to complete the proof. Since $aU_{1}=aV_{1}$, for each $h\in U_{1}$ there exists at least one $k\in V_{1}$ with $ah=ak$, and so each element $h\in U_{1}$ appears as a possible first co-ordinate of $B$, giving $p_{1}T_{1}=U_{1}$. Similarly, each element $k\in V_{1}$ appears as a possible second co-ordinate of $B$, giving $q_{n}T_{n}=V_{1}$. For $1\leq i\leq n-1$ we also have $y\in q_{i}T_{i}$ if and only if there exists $(h,k)\in B$ such that $y=q_{i}t_{i}(h,k)$ if and only if there exists $(h,k)\in B$ such that $y=p_{i+1}t_{i+1}(h,k)$ if and only if $y\in p_{i+1}T_{i+1}$. ∎
Putting Theorem 4.6 together with Theorem 4.21 yields the following:
Corollary 4.22.
Let $M$ be a strongly finitely right equated monoid. Then $\mathcal{S}(M)$ is weakly right coherent if and only if $M$ is strongly right ideal Howson.
Corollary 4.23.
If the monoid $M$ is left (right) LCM, then $\mathcal{S}(M)$ is weakly left (right) coherent.
Proof.
The result follows directly from Corollary 4.17 and Corollary 4.22. ∎
@@
Our next lemma will be used show that there exist finitely right equated monoids $M$ such that $\mathcal{S}(M)$ is not finitely right equated.
@@
In the proof of Lemma 4.24 we argued that, under the conditions of the Lemma, there exists a right annihilator congruence of the form $\mathbf{r}((\emptyset,x))$ that is not finitely generated for some $x\in M$. In fact, to determine whether $\mathcal{S}(M)$ is is finitely right equated, it always suffices to consider right annihilator congruences of the form $\mathbf{r}((\emptyset,x))$ for $x\in M$.
Proposition 4.26.
Let $M$ be a monoid. The expansion $\mathcal{S}(M)$ of $M$ is finitely right equated if and only if for all $a\in M$ the right annihilator congruence $\mathbf{r}((\emptyset,a))$ is finitely generated.
@@
The forward direction is trivial. Suppose then that each right congruence $\mathbf{r}((\emptyset,a))$ is finitely generated. Fix $(A,a)\in\mathcal{S}(M)$ and let $X$ be a finite generating set for $\mathbf{r}((\emptyset,a))$. It is easy to see that $\mathbf{r}((\emptyset,a))\subseteq\mathbf{r}((A,a))$: indeed, if $(\emptyset,a)(P,p)=(\emptyset,a)(Q,q)$, then $aP=aQ$ and $ap=aq$, from which it immediately follows that $A\cup aP=A\cup aQ$ and $ap=aq$, giving $(A,a)(P,p)=(A,a)(Q,q)$.
Now consider $((U,u),(V,v))\in\mathbf{r}((A,a))$. Arguing as in the proof of Theorem 4.21, setting $U=U_{1}\cup U_{2}$ and $V=V_{1}\cup V_{2}$ where $U_{1}=U\setminus a^{-1}A$, $U_{2}=U\cap a^{-1}A$, $V_{1}=V\setminus a^{-1}A$ and $V_{2}=V\cap a^{-1}A$, we find that
$(U,u)$ $=$ $(\emptyset,1)(U,u),\qquad(a^{-1}A,1)(U,u)=(a^{-1}A,1)(U_{1},u),\qquad(\emptyset,1)(U_{1},u)=(U_{1},u)$
$(V_{1},v)$ $=$ $(\emptyset,1)(V_{1},v),\qquad(a^{-1}A,1)(V_{1},v)=(a^{-1}A,1)(V,v),\qquad(\emptyset,1)(V,v)=(V,v),$
where $(A,a)(\emptyset,1)=(A,a)(a^{-1}A,1)$. Since there is an $X$-sequence from $U_{1}$ to $V_{1}$, setting $Y=X\cup\{((\emptyset,1),(a^{-1}A,1)),((a^{-1}A,1),(\emptyset,1))\}$ it follows that we can construct a $Y$-sequence from $U$ to $V$. Since $Y$ is finite, this completes the proof. ∎
Remark 4.27.
In light of Proposition 4.26 and the proof of Theorem 4.21, to determine whether $\mathcal{S}(M)$ is finitely right equated it suffices to focus on the right congruences $\mathbf{r}((\emptyset,a))$ such that $a\in M$ is not $\mathcal{L}^{*}$-related to an idempotent.
If $M$ is strongly finitely right equated, then we have shown in Theorem 4.21 that $\mathcal{S}(M)$ is strongly finitely right equated, that is, for any fixed $(A,a)\in\mathcal{S}(M)$ there is a finite set generating $W$ for $\mathbf{r}((A,a))$ with respect to which any two elements of $\mathbf{r}((A,a))$ are connected by a $W$-sequence with the same skeleton. Likewise, the proof of Theorem 4.15 shows that if $\mathcal{S}(M)$ is finitely left equated then for any $(A,a)\in\mathcal{S}(M)$ there is a finite generating set $Z$ for $\mathbf{l}((A,a))$ such that any pair of elements in $\mathbf{l}((A,a))$ may be connected via a $Z$-sequence having one of finitely many skeletons; this is equivalent to there being a bound on the length of $Z$-sequences required. Further, if $M$ is right cancellative, then we can take a single skeleton.
Question 4.28.
Characterise the monoids $M$ such that $\mathcal{S}(M)$ is finitely right equated. Does this characterisation imply that there is a bound on the length of sequences required?